OpenStudy (anonymous):

How do I factor completely the following polynomial; 396b^2-539c^2?

OpenStudy (amistre64):

with any luck its a difference of 2 squares

OpenStudy (amistre64):

\[a^2-b^2 = (a+b)(a-b)\]

OpenStudy (whpalmer4):

It is, but there's a constant factor to take out, too.

OpenStudy (whpalmer4):

Here's a hint: 360 * 11 = 396

OpenStudy (amistre64):

wheres my slide rule when i need it :)

OpenStudy (whpalmer4):

packed away in a box somewhere, just like mine!

OpenStudy (whpalmer4):

Uh, let's make that hint a little better: 36*11 = 396

OpenStudy (whpalmer4):

That means we need to factor \[11(36b^2-49c^2)\]which hopefully you recognize as a difference of two squares. Using @amistre64's formula, what are the factors?

OpenStudy (whpalmer4):

To avoid any letter confusion, I'll rewrite the formula as \[u^2-v^2 = (u+v)(u-v)\]\[36b^2-49c^2\]Comparing the formulas, we see that\[u^2=36b^2\]\[v^2=49c^2\]Can you solve those for \(u\) and \(v\)?

OpenStudy (anonymous):

Hey whpalmer, although ur answer is correct, is there a different form to enter 11(36b^2-49c^2)?

OpenStudy (whpalmer4):

we haven't factored it all the way yet...

OpenStudy (anonymous):

oh got it....

OpenStudy (whpalmer4):

so, if you figure out \(u\) and \(v\) in those little equations, the final answer will be \[11(u-v)(u+v)\]

OpenStudy (anonymous):

Thanks for ur assistance whpalmer, I appreciate it.